# BEVTraj: Map-Free End-to-End Trajectory Prediction in Bird’s-Eye View with Deformable Attention and Sparse Goal Proposals

### v1

@@ IV-E Ablation Study
We conducted an ablation study to evaluate the contribution of each module within BEVTraj to trajectory prediction performance. All experimental conditions follow the setup described in SectionIV-A, except for the inclusion or removal of the target module.
IV-E1 Pre-Encoder
We assessed the impact of the Pre-Encoder by comparing models with and without this component. As discussed in Section III-B, our model adopts a temporal bottleneck strategy, in which the temporal dimension is first compressed and later reintroduced during trajectory decoding. This design enables the model to reduce computational complexity during scene encoding while preserving essential temporal information for downstream prediction. Since both BEVTraj and MTR adopt temporal compression in their architectures (e.g., max pooling in MTR), they provide a meaningful basis for assessing the effectiveness of early temporal modeling prior to compression.
As shown in Table III, incorporating the Pre-Encoder improves all evaluation metrics except Miss Rate, with the most significant gain observed in minFDE. These findings indicate that applying temporal and social attention prior to temporal compression is effective in enhancing trajectory prediction performance. Specifically, temporal self-attention captures motion trends over time, while social attention models inter-agent interactions occurring within the same timestamp. The latter is particularly important, as it enables the model to capture fine-grained relational dependencies before temporal resolution is reduced.
This approach contrasts with modeling such interactions after temporal compression, where important short-term relational cues may have already been lost. By incorporating the Pre-Encoder, BEVTraj achieves a favorable balance between prediction accuracy and parameter efficiency, offering a design paradigm that may generalize to other models employing temporal bottleneck structures (e.g., MTR).
IV-E2 Sparse Goal Candidate Proposal
The SGCP module generates goal candidates that serve as initial endpoints for trajectory prediction and are iteratively refined through subsequent decoder layers. To assess its impact, we replaced SGCP with a variant that employs learnable queries without explicit conditioning on the target agent’s state or spatial context. As shown in Table III, incorporating SGCP consistently outperforms the baseline across all metrics, with the most significant improvement observed in minFDE, indicating enhanced accuracy in endpoint prediction within each mode.
This performance gain can be attributed to the SGCP’s architecture, which generates goal candidates based solely on the dynamic state of the target agent and static scene elements such as road topology—without leveraging inter-agent interactions or the scene context feature $\mathbf{C}$. Compared to full trajectory prediction, which requires reasoning over multi-agent dynamics and long temporal horizons, goal prediction is a simpler sub-task. Solving this sub-task early enables the model to encode the target agent’s intention into the content query $\mathbf{Q_{c}}$ as prior knowledge. Furthermore, the predicted goal candidates provide spatial anchors for the Initial Trajectory Prediction (ITP) module, helping downstream components align their predictions with more semantically plausible destinations.
TABLE III: Ablation study on the impact of different components.
@@ IV-E2 Sparse Goal Candidate Proposal
O O 1.4556 0.9438 8.4384 2.0527 0.3082
@@ IV-F Discussion
IV-F1 Effect of Map Range on Prediction Performance
As discussed in SectionIV-C, in the absence of pre-built HD maps, real-time HD map construction is inherently constrained by the sensor’s perceptual range. To examine the implications of this limitation, we analyze how the performance of HD map-based models varies with respect to map range, as illustrated in Figs. 8 and 9 and TablesIV and V. Here, the map range is defined relative to the ego vehicle’s position at the current timestamp. As the spatial range of the map increases, performance generally improves across most metrics for the majority of models, indicating that access to a wider spatial context enhances predictive capability.
To better assess the models’ ability to handle challenging prediction scenarios under different map ranges, we slightly modified the target agent selection criteria for the Argoverse 2 Sensor experiments. In particular, we excluded agents that are nearly stationary or moving in a straight line, as trajectory prediction becomes more difficult when agents move at higher speeds or make sharp turns. This adjustment reduces the number of easy cases in the evaluation set, resulting in overall higher error values in Table II compared to those reported in Table II, regardless of model type.
Although this trend generally holds, there are notable deviations, where a smaller map range yields better results or an extended range leads to degraded performance. These exceptions suggest that the optimal spatial extent of static map information may be influenced by model architecture or its capacity to encode contextual cues. Moreover, although this study employs ground-truth HD maps to isolate the effects of map range, real-world deployments must also account for the variability in the accuracy of real-time HD map construction, which may deteriorate as spatial coverage increases. These observations underscore the need for future work to jointly consider model-specific context integration capabilities and the reliability of map construction across varying spatial ranges.
The spatial constraints imposed by the sensor’s perceptual field apply not only to real-time HD map construction but also to BEVTraj. These findings indicate that BEVTraj performance may further improve with extended sensor range, as its deformable attention mechanism allows the model to selectively focus on informative regions within the BEV feature space. Future advancements in sensing technology or cooperative infrastructure—such as V2X communication—could expand the observable environment, thereby enhancing both perception and prediction capabilities.
Figure 8: Performance comparison of HD map-based models on the validation set of the nuScenes dataset. (a) minADE5. (b) minADE10. (c) minFDE1. (d) minFDE10. (e) Miss rate.
Figure 9: Performance comparison of HD map-based models on the validation set of the Argoverse 2 Sensor dataset. (a) minADE5. (b) minADE10. (c) minFDE1. (d) minFDE10. (e) Miss rate.
@@ IV-F1 Effect of Map Range on Prediction Performance
TABLE IV: Performance comparison of HD map-based models on the nuScenes validation set. The number in parentheses indicates the spatial range (in meters) of the input HD map (i.e., 30m, 50m, 100m).
@@
TABLE V: Performance comparison of HD map-based models on the Argoverse 2 Sensor validation set. The number in parentheses indicates the spatial range (in meters) of the input HD map (i.e., 30m, 50m, 100m).
@@
IV-F2 Ego-Vehicle Trajectory Prediction
In addition to forecasting the trajectories of surrounding agents, we also evaluated our model’s performance in predicting the ego vehicle’s future motion. Unlike conventional baseline models that rely solely on HD maps and historical trajectories, our approach leverages raw sensor inputs—such as LiDAR and camera data—to construct a richer and more adaptive representation of the scene.
As shown in Table VI, our model achieves comparable or slightly improved performance in ego trajectory prediction relative to baseline methods. This performance gain can be attributed to the model’s capacity to incorporate dynamic environmental cues and agent interactions, which are often underrepresented in HD map-based frameworks. Sensor-derived BEV representations offer real-time insights into nuanced scene dynamics and unexpected events within the ego vehicle’s perceptual field, thereby enhancing the model’s understanding of driving intent and scene context. This results in more accurate and responsive ego trajectory prediction.
Furthermore, the proposed model architecture is inherently extensible beyond trajectory prediction, offering a natural pathway toward motion planning applications. By leveraging real-time perception to infer feasible and context-aware future trajectories, the framework can be seamlessly integrated into downstream decision-making and path planning modules, underscoring its potential as a unified solution that bridges the gap between perception and planning in autonomous driving systems.
TABLE VI: Ego-vehicle trajectory forecasting results on the validation set of the nuScenes Dataset. BEVTraj achieves comparable performance to HD map-based baselines under the same spatial coverage constraint.
@@ V Conclusion
For future work, we aim to extend the framework by incorporating BEV feature-based object detection and tracking modules, enabling the joint processing of static and dynamic scene information within a unified representation space. Additionally, enhancing the temporal modeling capacity of BEV features through recurrent architectures or cross-sweep fusion techniques may further improve the model’s ability to capture long-term motion patterns. Investigating adaptive BEV grid structures could also help mitigate spatial limitations, allowing for more flexible and fine-grained trajectory prediction. We anticipate that these improvements will contribute to a more comprehensive and robust trajectory prediction system.
Furthermore, the techniques developed in this work are not limited to autonomous driving but can also be applied to broader domains, such as surveillance systems—where monitoring vehicle and human activities is critical for safety [38, 39, 40, 41, 42], —and robotics, where accurate trajectory prediction is essential for tasks such as human–robot interaction and navigation in dynamic environments [43].

### v2

@@ IV-E Ablation Study
We conduct comprehensive ablation studies to analyze the design of BEVTraj from three aspects: the contribution of core architectural components, the scalability and runtime trade-offs with respect to the number of modes, and key design choices in the decoder. Unless otherwise specified, all experimental settings follow Section IV-A.
TABLE V: Ablation study on the impact of different components.
@@
O O 1.3953 0.9051 8.1975 1.8964 0.2783
IV-E1 Component-wise Ablation
We first evaluate the contribution of the Pre-Encoder and SGCP. As shown in Table V, introducing the Pre-Encoder consistently improves overall accuracy—particularly minFDE—by modeling temporal dynamics and inter-agent interactions prior to temporal compression, while SGCP further enhances endpoint prediction by generating intention-aware goal candidates that provide reliable spatial anchors for downstream decoding. These results confirm that early temporal–social modeling and explicit goal proposal are both critical for robust trajectory forecasting.
IV-E2 Mode Scalability and Runtime Analysis
To validate that a small set of carefully selected, scene-adaptive goals can achieve strong performance without brute-force mode scaling, we analyze the effect of increasing the number of trajectory modes with learning-rate tuning and extended training. As shown in Table VI, adding more modes degrades most evaluation metrics without improving reliability. This behavior is consistent with prior findings: Shi et al. [46] demonstrate that when queries are not anchored to predefined intentions, increasing the number of modes results in unstable optimization and arbitrary mode switching. Taken together, these results indicate that a sparse set of scene-adaptive goals is both sufficient and more effective than brute-force mode scaling.
TABLE VI: Ablation study on the number of prediction modes.
Num Modes minADE5 minADE10 minFDE1 minFDE10 Miss Rate↓
10 (default) 1.3953 0.9051 8.1975 1.8964 0.2783
64 1.6070 1.0132 9.1218 2.1571 0.2667
Beyond accuracy, increasing the number of modes also reduces inference efficiency, as decoder runtime grows substantially with the number of modes (Table VII). Although deformable attention scales more favorably than global attention, naive mode scaling remains computationally inefficient. Overall, these results show that post-hoc selection from many modes is ill-suited for anchor-free designs such as BEVTraj, motivating SGCP to instead predict a small set of high-quality, scene-adaptive goals that better balance accuracy and efficiency.
TABLE VII: Runtime of the iterative trajectory decoder under different numbers of modes. Increasing the number of modes substantially increases latency, especially with vanilla attention.
Configuration Runtime (ms)
Mode 10 (deform attn) 19.30
Mode 10 (vanilla attn) 27.42
Mode 16 (deform attn) 19.35
Mode 64 (deform attn) 26.41
Mode 64 (vanilla attn) 85.85
Mode 128 (deform attn) 43.83
Mode 256 (deform attn) 83.64
IV-E3 Design Choices in the Decoder
We analyze key design choices in the decoder, focusing on the coordinate-system formulation for reference point prediction and the attention mechanism used for BEV feature aggregation in SGCP. As shown in Table VIII, using either ego-centric or target-centric coordinates alone leads to suboptimal performance: ego-centric modeling relies on unnormalized absolute positions, while target-centric modeling introduces structural misalignment with ego-centric BEV features. In contrast, jointly encoding both coordinate systems consistently achieves the best performance across all metrics, indicating that effective reference point prediction requires normalized relative motion cues together with spatial alignment to the BEV frame.
We further revisit the deformable attention formulation used in the decoder. Standard deformable attention does not explicitly encode positional information in keys, which limits its ability to reason about geometric constraints in BEV space. By explicitly constructing positional keys and decomposing attention into content and positional components, the decoder achieves lower displacement errors, as shown in Table IX. This result demonstrates improved spatial reasoning and goal localization, validating the effectiveness of explicit positional key modeling for trajectory prediction.
TABLE VIII: Ablation study on different coordinate-system configurations for decoder-side reference point prediction in SGCP. The baseline corresponds to the original SGCP design prior to the proposed revision.
Configuration minADE5 minADE10 minFDE1 minFDE10 Miss Rate↓
baseline 1.4556 0.9438 8.4384 2.0527 0.3082
ego-centric 1.3978 0.9056 8.4517 2.0267 0.2952
target-centric 1.4528 0.9444 8.0119 2.0069 0.2875
ec & tc 1.3953 0.9051 8.1975 1.8964 0.2783
TABLE IX: Ablation study on explicit positional key construction for deformable attention in the ITP/ITR stages.
Pos Key minADE5 minADE10 minFDE1 minFDE10 Miss Rate↓
X 1.4213 0.9377 9.8164 1.9349 0.2685
O 1.3953 0.9051 8.1975 1.8964 0.2783
@@ IV-F Discussion
While the previous sections focus on benchmark-level performance and component analysis, we further examine how BEVTraj behaves under practical deployment conditions. In this section, we analyze scene-level multi-agent consistency, sensing-range limitations, and ego-motion forecasting to better understand the system-level implications of the proposed design.
IV-F1 Multi-agent Trajectory Prediction
Method
To model the future behaviors of multiple interacting agents, we construct $S$ scene-level joint future predictions, each formed by selecting one of the modes per agent and combining them into a joint world [68]. A naive strategy is to independently choose the highest-probability mode for each agent. However, such independent selection over-relies on model confidence and often collapses the multi-modal hypothesis set, reducing diversity. Moreover, it fails to enforce cross-agent compatibility, potentially yielding interaction-infeasible joint futures. Therefore, we formulate joint-world construction as a constrained assignment problem with interaction-aware costs.
For each agent $i$, we assign its $K$ modes to the $S$ worlds by minimizing a weighted cost:
$\begin{split}C_{i}(s,k)=&\,w_{\mathrm{prob}}\,c_{\mathrm{prob}}(i,k)+w_{\mathrm{pair}}\,c_{\mathrm{pair}}(i,k;\mathcal{T}_{s})\\ &+w_{\mathrm{ego}}\,c_{\mathrm{ego}}(i,k)+w_{\mathrm{reg}}\,c_{\mathrm{reg}}(i,k),\end{split}$ (6)
where $c_{\mathrm{prob}}(i,k)=-\log p_{i,k}$ and $c_{\mathrm{pair}}(i,k;\mathcal{T}_{s})$ penalizes sustained close proximity to trajectories already cached in world $s$ (distance threshold with an $L$-consecutive criterion); $c_{\mathrm{ego}}$ is optional and $c_{\mathrm{reg}}$ is a lightweight motion regularizer.
We then solve the resulting minimum-cost one-to-one assignment using the Hungarian algorithm, with $\mathcal{T}_{s}$ updated sequentially so that $c_{\mathrm{pair}}$ accounts for agents already assigned to each world. This produces diverse joint worlds that better respect interaction feasibility for system-level evaluation.
Figure 14: Visualization of BEVTraj’s multi-agent predictions for several agents that exhibit strong interaction relevance with the ego-vehicle. Red indicates the target agents, and green indicates the ego vehicle. For compact presentation, multi-worlds predictions are aggregated into a single visualization, and only the top three most probable modes are shown.
System-level Multi-agent Evaluation
We run inference for all agents in a scene, assemble them into S=6S{=}6 joint multi-worlds using our matching procedure, and evaluate them as $K$ scene-level predictions under the Argoverse 2 [69] protocol using its standard system-level metrics (avgMinADE, avgMinFDE, actorMR, and ActorCR). In addition, we introduce NodeCR, a node-level collision rate that measures the fraction of actor–timestep nodes whose pairwise distance violates a threshold for at least $L$ consecutive timesteps (default $d_{\mathrm{th}}{=}1.6\,\mathrm{m},\,L{=}2$), capturing sustained overlap risk.
Analysis
Fig. 14 shows qualitative examples of BEVTraj’s multi-agent trajectory predictions. Although BEVTraj operates without HD-map priors, the proposed Matching post-processing substantially improves the safety of the assembled joint futures, yielding the lowest collision-related rates after Matching (NodeCR=0.0000=0.0000, ActorCR=0.0002=0.0002) and the lowest miss rate (ActorMR=0.4178=0.4178) in Table X, outperforming HD-map-based baselines. Here, ActorCR counts an actor as collided if its minimum distance to any other actor falls below $1.0\,\mathrm{m}$ at any future timestep, while NodeCR measures sustained overlap risk as the fraction of actor–timestep nodes that violate a distance threshold for at least $L$ consecutive timesteps (default $d_{\mathrm{th}}{=}1.6\,\mathrm{m},\,L{=}2$).
TABLE X: Performance comparison under the multi-world prediction evaluation protocol on the nuScenes validation set (lower is better).
[…]
Autobot 2.6421 6.4028 0.7648 0.0002 0.0022
MTR 2.1352 5.4476 0.7229 0.0005 0.0030
Wayformer 2.1675 5.7544 0.7537 0.0001 0.0018
DEMO 1.8843 4.8902 0.6864 0.0002 0.0025
BEVTraj (Ours) 1.9285 4.9791 0.4178 0.0000 0.0002
IV-F2 Effect of Map Range on Prediction Performance
As discussed in Section IV-C, real-time HD map construction is inherently limited by the sensor’s perceptual range. To analyze how this constraint affects prediction performance, we evaluated HD map-based models under varying map ranges (Tables XI and XII). Across most models and metrics, expanding the spatial range improves performance, confirming that broader contextual information benefits trajectory prediction. However, the gains are not uniform across architectures, and some models fail to effectively exploit extended context, occasionally showing limited or degraded performance. This observation highlights that the usefulness of additional spatial information depends on how well a model integrates contextual cues. In contrast, BEVTraj is explicitly designed to aggregate informative regions through deformable attention, suggesting that it can more effectively capitalize on expanded sensing coverage. Consequently, extending the perceptual range—through improved sensors or cooperative infrastructure such as V2X—may further enhance its prediction accuracy.
@@ IV-F2 Effect of Map Range on Prediction Performance
TABLE XI: Performance comparison of HD map-based models on the nuScenes validation set. The number in parentheses indicates the spatial range (in meters) of the input HD map (i.e., 30m, 50m, 100m).
@@
TABLE XII: Performance comparison of HD map-based models on the Argoverse 2 Sensor validation set. The number in parentheses indicates the spatial range (in meters) of the input HD map (i.e., 30m, 50m, 100m).
@@
IV-F3 Ego-Vehicle Trajectory Prediction
In addition to predicting surrounding agents, we evaluated the model’s ability to forecast the ego vehicle’s future trajectory. As shown in Table XIII, our method achieves comparable or slightly improved performance over HD map-based baselines. Notably, ego trajectory prediction tends to be more reliable than surrounding-agent prediction. This advantage stems from the use of onboard sensor data, which provides rich, real-time cues that are directly coupled with the ego vehicle’s own motion and immediate environment. Such ego-centric observations offer stronger behavioral signals than those available for other agents, enabling more accurate and stable ego trajectory forecasting.
Furthermore, although BEVTraj is primarily designed for trajectory prediction, its sensor-driven BEV representation and interaction-aware modeling are conceptually aligned with the requirements of motion planning. These properties suggest that the proposed framework can offer useful insights and transferable components for planning research, supporting closer integration between perception, prediction, and decision-making.
TABLE XIII: Ego-vehicle trajectory forecasting results on the validation set of the nuScenes Dataset. BEVTraj achieves comparable performance to HD map-based baselines under the same spatial coverage constraint.
@@ V Conclusion
Beyond autonomous driving, the techniques developed in this work are not limited to a single application domain but can also be extended to broader areas, such as surveillance systems—where monitoring vehicle and human activities is critical for safety [70, 71, 72, 73, 74, 75]—and robotics, where accurate trajectory prediction is essential for tasks such as human–robot interaction and navigation in dynamic environments [76].